# RefVTON: person-to-person Try on with Additional Unpaired Visual Reference

### v3

@@ Abstract
Figure 1: In-the-wild try-on results generated by our RefVTON model with a p2p style, trained on person and garment images from our Virtual Fitting with Reference (VFR) dataset.The first row demonstrates our mask-free try-on capability, where the garment is transferred directly to the target person without masks or pose estimation. The second row shows our additional-reference try-on mode, in which extra visual references are incorporated to enhance structural accuracy, texture fidelity, and overall realism.
Figure 2: The effect of using reference images for the virtual try-on task. From left to right in the three middle subfigures are: (i) results generated without using reference images during either training or inference; (ii) results generated by a model trained and inferred with reference images. Incorporating reference images consistently improves the try-on quality and authenticity in both training and inference stages. Please zoom in for more details.
@@ 1 Introduction
Based on the above observation, we propose RefTON, a flux-based person-to-person virtual try-on framework that achieves strong performance without relying on any external models or auxiliary components, while being further enhanced by additional reference images that offer more accurate and context-aware guidance for the try-on model. First, to ensure the best performance, we adopt the powerful image editing model flux-kontext as our base and apply adaptation on its position index for RoPE [su2024roformer] to make it suitable for muti-conditonal virtual try-on tasks. Similar to [chong2024catvton], RefTON eliminates the need for auxiliary inputs such as segmentation masks using a two-stage training strategy, allowing for simple inference with only a source image and the target garment as inputs. Second, we introduce the use of images of a different person clothed in the target garment as the visual references, like the model images in online shopping, which better reflect users’ real-world behavior when choosing clothes and enable the preservation of fine garment details that existing methods cannot achieve. To achieve these objectives, we propose a reference data generation pipeline, by which construct a dataset with supplementary reference images and use unpaired person-cloth sample to train our own model to utilize reference image as additional visual guidance. These improvements empower RefTON to achieve both a simplified model structure and a streamlined inference process, while simultaneously delivering superior generation results. To the best of our knowledge, RefTON is the first to support both mask-based and mask-free try-on within a single model while incorporating reference images into the pipeline.
@@
  • We designed an reference data generation framework to create the necessary reference images for both the clothing and target ground truth samples. Based on this pipeline, we built the VFR dataset upon existing benchmarks (e.g., VITON-HD, DressCode, ViViD), providing a robust new resource to improve the practicality and evaluation of virtual try-on models.
  • We present an adaptation of the Flux-Kontext I2I model with a modified positional indexing mechanism for handling multi-conditional/resolution inputs in virtual try-on. Leveraging this architecture and a two-stage training strategy, our model effectively supports both mask-based and person-to-person try-on tasks, achieving state-of-the-art results and strong generalization to in-the-wild person–clothing images.
@@ 2 Related Works
Generative modeling has advanced rapidly, with diffusion models (DMs) [sohl2015deep], score-based generative models (SGMs) [song2020score], and flow-based methods emerging as leading paradigms. Flow-based methods [dinh2016density, dinh2014nice, kingma2018glow] have evolved to address inefficiencies in traditional continuous normalizing flows (CNFs)—which require expensive backpropagation through ODE solvers during training [Chen2018NeuralOD]. A breakthrough in this area is Flow Matching (FM) [Lipman2022FlowMF], which learns a time-dependent vector field to deterministically transport a simple prior distribution to the target data distribution. Unlike CNFs, FM uses a simulation-free training objective that avoids numerical integration at training time, significantly reducing computational overhead. This approach not only achieves comparable or superior sample quality to DMs but does so with far fewer sampling steps, as it directly parameterizes the probability flow instead of learning a stochastic reverse process [Lipman2022FlowMF]. The Flux-Kontext architecture offers a powerful and flexible method for image editing. Input images are encoded into latent embeddings, which are then flattened into sequences and concatenated with Gaussian noise $\bm{\epsilon}$. Our model builds upon this architecture, leveraging its flexibility and ability to manipulate and edit images in a controlled and simplified way.
2.1 Diffusion-based Virtual Tryon
Diffusion models have advanced rapidly in recent years [Ho2020DenoisingDP, Rombach2021HighResolutionIS], leading to a wide range of diffusion-based approaches in the virtual try-on domain [kim2024stableviton, zhu2023tryondiffusion, choi2021viton]. Stable Diffusion [morelli2023ladi, kim2024stableviton], with its flexible inpainting and text-guided capabilities, has become widely adopted for virtual try-on tasks. DiffusionCLIP [kim2022diffusionclip] further incorporates CLIP loss to refine the generated images. DCI-VTON [gou2023taming] follows a traditional two-stage pipeline, first warping clothing to align with the body and then fusing the warped garment with the person’s image. Subsequently, IDM-VTON [choi2021viton] introduces a dedicated GarmentNet module that guides the garment structure and appearance, further improving visual realism. CatVTON [chong2024catvton], OmniTry [feng2025omnitry], Any2AnyTryon [guo2025any2anytryon], and OmniVTON [yang2025omnivton] investigate p2p virtual try-on methods that enable mask-free try-on. However, they either require additional conditions, such as human pose, or cannot handle mask-based and mask-free tasks simultaneously. Despite these advances, diffusion-based virtual try-on methods still face notable limitations in practical application and technical completeness, leaving room for further research and optimization.
@@ 2.1 Diffusion-based Virtual Tryon
In addition, p2p tryon methods such as TryOffDiff [velioglu2025enhancing] and ViTON-GUN [ViTON-GUN] rely on a “try-off–then–try-on” pipeline, which could easily lose garment details and accumulate errors. In contrast, our RefVTON skips the try-off stage, directly fitting the target garment onto the person. Guided by the reference image, this approach ensures superior preservation of both garment structure and material fidelity. In summary, existing diffusion-based virtual try-on methods either (1) rely on excessive auxiliary annotations or (2) lack support for reference images of clothed individuals. Our RefTON adapts the virtual try-on task to the Flux-Kontext model to address these limitations, combining person-to-person, reference-aided diffusion training while supporting both mask-based inpainting and mask-free editing for virtual try-on.
@@
Figure 3: The pipeline of our two-stage training strategy: (a) In the first stage, which follows a similar paradigm to mask-based try-on approaches, the model is trained on masked person images to generate person images wearing random garments for the next stage training. (b) In the second stage, the synthesized person images produced in the first stage, along with the target garment and additional reference images (optional), are jointly used as inputs to train a person-to-person virtual try-on model that directly fits the target cloth onto the person’s body.
@@ 3.1 Preliminary
We consider the problem of generating images conditioned on an embedding $\bm{y}$, which may encode garment images, semantic maps, human pose, or other modality-specific control signals. Let $\bm{x}$ denote the latent image representation obtained from a VAE encoder. The goal of flux.1 [flux2024, labs2025flux1kontextflowmatching] is to approximate the conditional distribution $p(\bm{x}\mid\bm{y})$ by learning a time-dependent velocity field $\bm{v}(\bm{x},\bm{y},t)$ that transports a sample from a simple prior $p_{0}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ at $t=0$ to the data distribution $p_{\text{data}}(\bm{x}|\bm{y})$ at $t=1$. The dynamics of the conditional probability density $p(\bm{x},t|\bm{y})$ over time $t$ are governed by the continuity equation:
@@
In the virtual try-on setting, let $\bm{x}_{i}$ denote the image of a person wearing the target cloth, and let $\bm{y}_{i}$ represent a collection of conditional inputs, including the cloth-agnostic image $\bm{a}_{i}$, the target cloth $\bm{c}_{i}$, the target person images $\bm{p}_{i}$, the visual references $\bm{r}_{i}$, and others. Formally, we write $\bm{y}_{i}=[\bm{a}_{i},\bm{c}_{i},\bm{d}_{i},\dots].$ The objective is to progressively transform a Gaussian noise sample $\bm{\epsilon}$ into the target image $\bm{x}_{i}$ guided by conditions $\bm{y}_{i}$.
@@ 3.2 Person To Person Virtual Try-on Model with Two Stage Training
Unlike prior methods that first mask the target clothing region to create an agnostic image, our goal is to fit the target garment onto the person directly, without relying on additional conditions such as densepose [guler2018densepose] or segmentation masks. To achieve this, we train a diffusion model on pairs of clothing images $\bm{c}_{i}$ and person images wearing different garments $\bm{\bar{p}}_{i,\bm{c}_{j}}$, as shown in Fig. 3(b). This requires constructing a training set of unpaired images of the same person in different clothes, denoted as $\bm{\bar{p}}_{i,c_{j}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$, where the subscript $c_{j}$ indicates a garment other than $\bm{c}_{i}$.
@@
Unfortunately, such unpaired cloth–person data pairs $\bm{\bar{p}}_{i,c_{j}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$ are not available in current open-source benchmarks and most of them only contain data pair $\bm{c}_{i},\bm{p}_{i,c_{i}}]$. Therefore, we synthesize unpaired person images $\bm{\bar{p}}_{i,c_{j}}$ using a mask-based try-on model, which serves as the first stage of our RefTON training strategy. As illustrated in Fig. 3(a), we train a virtual try-on model using agnostic person images $\bm{a}_{i}$, clothing images $\bm{c}_{i}$, densepose maps $\bm{d}_{i}$, warp masks $\bm{m}i$, and other auxiliary conditions as inputs. During inference, a random unpaired garment $\bm{c}_{j}$ is selected to generate the corresponding synthesized person image $\bm{\bar{p}}_{i,c_{j}}$. To ensure the quality of the synthesized person image $\bm{\bar{p}}_{i,c_{j}}$, the agnostic–cloth pairs $\bm{a}_{i},\bm{c}_{i}]$ used for training must belong to the same garment category (e.g., if $\bm{a}_{i}$ is from the “dresses” subset, the selected garment $\bm{c}_{j}$ should also come from “dresses” rather than “upper body” or “lower body”). Otherwise, the generated person image may appear unrealistic due to mismatches between the clothing mask region and the target garment (e.g., fitting a skirt onto the upper body or a shirt onto the lower body).
Figure 4: Adaptation of a three-channel position index: the first channel encodes different conditional inputs, while the second and third channels provide spatial positional information for adapting the resolution of the target inputs.
3.3 Multi-input Adaptation of RefTON
After obtaining the unpaired person images, we train our person-to-person model by replacing the agnostic images $\bm{a}_{i}$ with the corresponding person images $\bm{p}_{i}$, which the first stage model generates. To enable the model to handle both masked and unmasked inputs flexibly, we randomly feed agnostic and person images with equal probability (50%) during training. Instead of training a try-on model from scratch, we adopt Flux-Kontext [labs2025flux1kontextflowmatching] as the foundation model, freeze the parameters of its encoder and decoder, and apply Low-Rank Adaptation (LoRA) [hu2022lora] to the MM-DiT blocks for parameter-efficient fine-tuning. The training objective follows the same flow-matching loss as defined in Equation 2, where $\bm{a}_{i}$ and $\bm{p}_{i}$ are alternately used to ensure compatibility with both input types. To enable the model to extract visual information from the target garment worn by another person, an additional reference image $\bm{r}_{i}$ is provided as input with a certain probability ($25\%$). Its latent embeddings are appended to the latent feature sequence $\bm{\epsilon}_{i},\bm{a}_{i},\bm{c}_{i}]$ after the image encoder, as illustrated in Fig. 4.
Some prior work [guo2025any2anytryon] leverages adaptive position embeddings for virtual try-on. Similarly, Flux-Kontext employs an index generator to produce a position index of shape $[L,3]$ for the image embedding sequence, which is then transformed into Rotary position Embeddings (RoPE) to provide spatial cues. The first channel distinguishes Gaussian noise $\bm{\epsilon}_{i}$ from conditional embeddings, while the second and third encode horizontal and vertical coordinates (Fig. 4).
Building on this design, our model converts conditional inputs into latent feature after Auto-Encoder and flatten into $L$ $D$-dimensional latent embeddings, denoted as $\bm{a}_{i},\bm{c}_{i},\bm{r}_{i}]$, and concatenates them with Gaussian noise embeddings $\bm{\epsilon}_{i}$ along the sequence dimension. This formulation allows conditional inputs of multi-resolutions to be concatenated—unlike [guo2025any2anytryon], which concatenates noisy and conditional inputs directly in pixel space, requiring them to share the same image size. Each embedding is assigned a three-channel position index: the first channel is extended from a binary flag to four discrete values to distinguish multiple conditions, and the second and third channels store integer spatial coordinates. To handle multi-resolution conditions, we further rescale these position indices by the resolution ratio between the target and conditional images, ensuring numerical consistency across resolutions. The modified DiT position index design and the overall RefTON architecture are illustrated in Fig. 4.
Figure 5: The overall pipeline of generating the reference images. We first generate the appearance descriptions using Qwen2.5-VL [bai2025qwen2], and then concatenate the appearance with the corresponding actions and outfits to construct the positive and negative prompts, as shown in (a). Subsequently, the images and the textual prompts are fed into the Editing Model, which generates photos of individuals wearing the same clothes. These results are reference images for each image–cloth pair, as shown in (b).

### v5

# RefTon: Reference person shot assist virtual Try-on

@@ Abstract
Figure 1: In-the-wild try-on results produced by our RefTON model. The first row demonstrates our mask-free try-on capability, where the garment is transferred directly to the target person. The second row shows our additional-reference try-on mode, in which extra visual references are incorporated to enhance structural accuracy, texture fidelity, and overall realism.
11footnotetext: Equal contribution.22footnotetext: Corresponding author
@@ 1 Introduction
Figure 2: Effect of reference images in virtual try-on. From left to right: inputs (agnostic, cloth, reference), results without reference images, results with reference images, and ground truth. Using reference images in both training and inference improves visual fidelity and preserves fine details (e.g., transparent materials and lace collars). Zoom in for a better comparison.
Based on the above observation, we propose RefTon, a flux-based person-to-person virtual try-on framework that achieves strong performance without relying on any external models or auxiliary components, while being further enhanced by additional reference images that offer more accurate and context-aware guidance for the try-on model. First, to ensure the best performance, we adopt the powerful image editing model Flux-kontext as our base and apply adaptation on its position index for RoPE [su2024roformer] to make it suitable for multi-condition/resolution virtual try-on inputs. Similar to [chong2024catvton], RefTon eliminates the need for auxiliary inputs such as segmentation masks using a two-stage training strategy, allowing for simple inference with only a source image and the target garment as inputs. Second, we introduce the use of images of a different person clothed in the target garment as the visual references, like the model images in online shopping, which better reflect users’ real-world behavior when choosing clothes and enable the preservation of fine garment details that existing methods cannot achieve. To achieve these objectives, we propose a reference data generation pipeline, by which we construct a dataset with supplementary reference images and use unpaired person-cloth samples to train our own model to utilize reference images as additional visual guidance. These improvements empower RefTon to achieve both a simplified model structure and a streamlined inference process, while simultaneously delivering superior generation results.
@@
  • We designed a reference data generation framework to create the necessary reference images for both the clothing and target ground truth samples. Based on this pipeline, we built the VFR dataset upon existing benchmarks (e.g., VITON-HD, DressCode, ViViD), providing a robust new resource to improve the practicality and evaluation of virtual try-on models.
  • We present an adaptation of the Flux-Kontext I2I model with a modified Rescaled Position Indexing mechanism to support flexible multi-conditional and multi-resolution inputs, along with a two-stage training strategy for virtual try-on. Our framework enables integration of varying numbers and types of reference images and effectively supports mask-based and person-to-person try-on within a single model. It achieves state-of-the-art performance and demonstrates strong generalization to in-the-wild person–clothing scenarios.
@@ 2.1 Generative Model via Flow Matching
Generative modeling has rapidly progressed with diffusion models (DMs) [sohl2015deep], score-based generative models (SGMs) [song2020score], and flow-based methods. Recent work has explored controllable generation, layout-conditioned synthesis, unified planning-generation frameworks [cheng2024hico, he2025plangen], as well as diffusion transformer variants for multilingual generation, efficient sampling, and fine-grained control [Liu_2025, ma2025nami, liu2025nanocontrollightweightframeworkprecise, gong2025cta]. Flow-based methods [dinh2016density, dinh2014nice, kingma2018glow] have been further advanced to address the inefficiency of continuous normalizing flows (CNFs), which require costly backpropagation through ODE solvers during training [Chen2018NeuralOD]. Flow Matching (FM) [Lipman2022FlowMF] mitigates this limitation by learning a time-dependent vector field that deterministically transports a simple before the data distribution, using a simulation-free objective that avoids numerical integration during training. By directly parameterizing the probability flow, FM achieves competitive or superior sample quality compared to diffusion models with significantly fewer sampling steps. Our method builds upon the Flux-Kontext architecture, where input images are encoded into latent representations, flattened into sequences, and concatenated with Gaussian noise $\bm{\epsilon}$. It is also closely related to recent FLUX-based appearance editing and transfer approaches [zhu2025flux].
2.2 Diffusion-based Virtual Try-on
Diffusion models [Ho2020DenoisingDP, Rombach2021HighResolutionIS] have enabled significant progress in garment or makeup transfer [kim2024stableviton, zhu2023tryondiffusion, choi2021viton, zhu2025flux]. Leveraging the flexibility of Stable Diffusion [morelli2023ladi, kim2024stableviton], prior works exploit text guidance and inpainting for garment synthesis. Extensions such as DiffusionCLIP [kim2022diffusionclip] introduce semantic control via CLIP, while methods like DCI-VTON [gou2023taming] and IDM-VTON [choi2021viton] adopt two-stage pipelines to align and fuse garments, improving structural consistency.
@@ 2.2 Diffusion-based Virtual Try-on
Recent approaches, including CatVTON [chong2024catvton], OmniTry [feng2025omnitry], Any2AnyTryon [guo2025any2anytryon], and OmniVTON [yang2025omnivton], explore person-to-person try-on without explicit masks. However, these methods typically rely on additional conditions (e.g., pose) or fail to unify mask-based and mask-free settings. Moreover, P2P pipelines such as TryOffDiff [velioglu2025enhancing] and ViTON-GUN [ViTON-GUN] adopt a “try-off–then–try-on” strategy, which introduces error accumulation and loss of garment details. In contrast, our method directly leverages the reference image, avoiding the try-off stage and better preserving garment structure and material fidelity. In summary, existing diffusion-based virtual try-on methods either rely on heavy auxiliary annotations or lack support for clothed reference images. To address these limitations, we propose RefTON, which adapts the virtual try-on task to the Flux-Kontext framework, enabling end-to-end, reference-guided generation while supporting both mask-based inpainting and mask-free editing.
@@
Figure 3: Overview of our two-stage training strategy. (a) Stage 1 follows mask-based try-on paradigms to generate person images wearing random garments from masked inputs, providing training data for Stage 2. (b) Stage 2 uses synthesized person images, along with the target garment and optional reference images, to train a person-to-person virtual try-on model that directly fits the garment onto the person.
@@ 3.1 Preliminary
We consider the problem of generating images under the condition $\bm{y}$, which may represent garment images, semantic maps, human pose, or other modality-specific control signals. Let $\bm{x}$ denote the latent image representation obtained from a VAE encoder. The goal of Flux.1 [flux2024, labs2025flux1kontextflowmatching] is to approximate the conditional distribution $p(\bm{x}\mid\bm{y})$ by learning a time-dependent velocity field $\bm{v}(\bm{x},\bm{y},t)$ that transports a sample from a simple prior $\mathcal{N}(\bm{x};\bm{0},\bm{I})$ at $t=0$ to the data distribution $p_{\text{data}}(\bm{x}|\bm{y})$ at $t=1$. The dynamics of the conditional probability density $p(\bm{x}|\bm{y},t)$ over time $t$ are governed by the continuity equation:
@@
In the virtual try-on setting, let $\bm{x}_{i}$ denote the image of a person wearing the target cloth, and let $\bm{y}_{i}$ represent a collection of conditional inputs, including the cloth-agnostic image $\bm{a}_{i}$, the target cloth $\bm{c}_{i}$, the visual references $\bm{r}_{i}$, and others. Formally, we write $\bm{y}_{i}=[\bm{a}_{i},\bm{c}_{i},\dots].$ The objective is to progressively transform a Gaussian noise sample $\bm{\epsilon}$ into the target image $\bm{x}_{i}$ guided by conditions $\bm{y}_{i}$.
@@ 3.2 Person To Person Virtual Try-on Model with Two Stage Training
Our goal is to dress the person directly with the target garment, without relying on auxiliary conditions such as DensePose [guler2018densepose] or segmentation masks. To achieve this, we train a diffusion model on clothing-person pairs $(\bm{c}_{i},\bm{\bar{p}}_{i,\bm{c}_{j}})$ to generate the target image $\bm{p}_{i,\bm{c}_{i}}$, where the person wears the target garment, as shown in Fig. 3(b). This setup requires unpaired triplets $\bm{\bar{p}}_{i,\bm{c}_{j}},\bm{c}_{i},\bm{p}_{i,\bm{c}_{i}}]$, where $\bm{c}_{j}\neq\bm{c}_{i}$. However, existing open-source benchmarks typically provide only paired data $\bm{c}_{i},\bm{p}_{i,\bm{c}_{i}}]$. Following the two-stage training strategy of CATVTON [chong2024catvton], we adopt a similar pipeline for our try-on model and further exploit richer conditions, including agnostic masks and DensePose, to improve unpaired image generation. Specifically, in the first stage of RefTon training, we synthesize unpaired person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$ using a mask-based try-on model.
@@
As illustrated in Fig. 3(a), we train a virtual try-on model using agnostic person images $\bm{a}_{i}$, clothing images $\bm{c}_{i}$, densepose maps $\bm{d}_{i}$, warp masks $\bm{m}_{i}$ as inputs. During inference, a random unpaired garment $\bm{c}_{j}$ is selected to generate the corresponding synthesized person image $\bm{\bar{p}}_{i,c_{j}}$. To ensure the quality of the synthesized person image $\bm{\bar{p}}_{i,c_{j}}$, the agnostic–cloth pairs $\bm{a}_{i},\bm{c}_{i}]$ used for training must belong to the same garment category (e.g., if $\bm{a}_{i}$ is from the “dresses” subset, the selected garment $\bm{c}_{j}$ should also come from “dresses” rather than “upper body” or “lower body”). Otherwise, the generated person image may appear unrealistic due to mismatches between the clothing mask region and the target garment (e.g., fitting a skirt onto the upper body or a shirt onto the lower body).
After obtaining the unpaired person images, we train the person-to-person model with either the agnostic image $\bm{a}_{i}$ or the generated person image $\bm{\bar{p}}_{i,c_{j}}$ from the first stage, sampled with equal $50\%$ probability. Instead of training a try-on model from scratch, we freeze the encoder and decoder of Flux-kontext [labs2025flux1kontextflowmatching] and fine-tune only the transformer blocks using Low-Rank Adaptation [hu2022lora]. The model is optimized with the flow-matching objective in equation. 2. In addition, to help the model capture garment appearance from another person, we provide an extra reference image $\bm{r}_{i}$ with a probability of $25\%$, as illustrated in Fig. 3(b).
Figure 4: Rescaled three-channel position index. (a) The first channel encodes conditional inputs, while the second and third capture spatial positions for varying resolutions. (b) The concatenated indices are used to generate RoPERoPE, which is applied to the latent sequence in the attention mechanism.
3.3 Multi-input Training and Adaptation
The latent embeddings are concatenated into the sequence $\bm{\epsilon}_{i},\bm{a}_{i},\bm{c}_{i}]$ after image encoding in vanilla Flux-kontext, and a index generator produces a position index of shape $[L,3]$, which is further converted into Rotary Position Embeddings (RoPE) for self-attention. Its first channel is a binary mask that distinguishes Gaussian noise $\bm{\epsilon}_{i}$ from conditional image/text inputs, while the second and third channels encode horizontal and vertical coordinates, respectively (Fig. 4). However, this binary design is insufficient for our setting, where the model must handle multiple heterogeneous image conditions. Therefore, we extend the first channel from a binary flag to discrete condition labels, allowing the transformer to distinguish different input types such as person, garment, and reference images.
Specifically, our model follows the Flux-kontext to encode each image condition independently into latent patches, and concatenates to form a sequence such as $\bm{a}_{i},\bm{c}_{i},\bm{r}_{i},\ldots]$, which is then concatenated with the noise latent $\bm{\epsilon}_{i}$ along the sequence dimension. For each condition, we generate an individual three-channel position index: the first channel indicates the condition identity, and the second and third channels store integer spatial coordinates rescaled by the resolution ratio between the target image and the corresponding condition image, which preserves spatial alignment across different resolutions. This Rescaled Position Index design enables flexible integration of multi-type and multi-resolution image conditions within a unified DiT framework. Similar indexing strategies have been explored in prior work [guo2025any2anytryon], where positional indices are constructed on a concatenated image canvas, placing different conditions in a shared spatial layout. In contrast, our method assigns positional indices to each condition independently rather than generating them from a pixel-space concatenated canvas, resulting in a more flexible formulation for multi-resolution inputs. The modified DiT position indexing scheme is illustrated in Fig. 4.
Figure 5: Reference image generation pipeline. Given a target image of a European woman wearing the target upper-body garment, Qwen2.5-VL [bai2025qwen2] produces an appearance description (e.g., ”a European woman with brown straight hair”) and an opposite description (e.g., ”an African woman with curly blonde hair”). In (a), the opposite description is combined with the action and non-target clothing attributes to form the positive prompt, while the appearance description is used as the negative prompt. In (b), the image and prompts are fed into Flux-kontext to generate reference images of different individuals wearing the target garment.